OpenStudy (anonymous):

The price of a cup of coffee was $2.30 yesterday. Today, the price rose to $2.55 . Find the percentage increase. Round your answer to the nearest tenth of a percent.

OpenStudy (anonymous):

2.30 * x = 2.55, where x = the increase plus 100%. Rearranged, x = 2.55/2.3 = 1.11 (rounded to 2 decimal places) So 2.55 is 1.11*2.30, i.e. 111% of 2.30. The increase is 11% :)

OpenStudy (anonymous):

Woops, I misread your question! x = 2.55/2.3 = 1.109 (3 decimal places) Increase is 10.9% (to the nearest 10th of a percent)